# Project page: https://xiaohanlei.github.io/projects/GaussNav/

### v2

@@ 4.4 Efficiency Analysis
By applying our grouping strategy based on semantic labeling, the search space is considerably narrowed. We count the number of different categories of object instances within the CrMo8WxCyVb-floor--2.9 scene, as demonstrated in Tab. 3. To locate a “chair” — assuming we render each object instance from three unique viewpoints — the resulting search space is reduced to merely $3\times 11=33$ observations. We render three observations of a single instance to ensure rendering quality, which is achievable only when the new viewpoint largely overlaps with the training views. This optimization yields a significant improvement in time efficiency.
4.5 Error Analysis
The performance of the proposed model is still far from perfect. We would like to understand the error modes for future improvements. Our analysis identifies two sources of error: the first being the model’s inability to consistently match the target from instance renderings, and the second, inaccuracies in goal localization. To quantify the impact of these error sources, we conduct an evaluation of our model using a ground truth Match module and an accurate goal localization. The first one means agent can correctly recognize the target from candidate observations, and the second suggests agent is directly provided with the ground truth goal position. The variant equipped with a ground truth Match module (GaussNav w. GT Match in Tab. 2) shows that Success can be enhanced by approximately 0.127 (rows 5 and 6 in Tab. 2). Furthermore, when the model is augmented with both ground truth Match and Goal Localization, denoted as GaussNav w. GT Goal Localization, we observe an increase in Success from 0.850 to 0.946, as indicated in rows 6 and 7 in Tab. 2. Improvements in the first error source may be achievable through the development of a more robust re-identification algorithm. As for the second source, a more precise Grounding strategy could yield better results. These insights not only highlight the model’s current shortcomings but also chart a course for subsequent refinement efforts.
@@ 4.6 Gaussian Construction Results
In Fig. 4, we present an evaluation of the rendering quality produced by our Semantic Gaussian Construction method on the HM3D validation dataset [41]. To align with the constraints of IIN task [19], we divide each scene within HM3D into separate floors and restrict the agent’s movement to within a single floor, as the IIN task [19] inherently ensures that both the agent’s starting location and the target’s position are on the same floor.
@@
To quantitatively analyze the results in Fig. 4, we observe that the rendering results exhibit a bifurcated trend. For instance, in scenes with indices 29 and 30, the rendered images achieve a high PSNR of up to 40 and a near-zero depth rendering error. However, the rendering performance for scene 10 is suboptimal. We hypothesize that this polarized rendering quality across different scenes can be attributed to the discrepancy between the simulation and reality. This is evident in Fig. 5, where some renderings from the HM3D dataset using Habitat simulator exhibit low fidelity, particularly in highly textured environments. High-quality reconstruction in such intricate settings is difficult, and utilizing suboptimal renderings as a basis for 3D environment reconstruction can further degrade the quality of the final output. In light of this, for scenes that are poorly reconstructed, we maintain consistency by using the original training views, rather than attempting to render novel views which would likely result in a diminished quality.
5 Conclusion
@@
In this work, we introduce a modular approach for visual navigation, i.e., Gaussian Splatting for Visual Navigation (GaussNav). Previous map-based methods largely focus on building 2D BEV map, which cannot represent the 3D geometry and detailed features in a scene. To this end, we propose a novel map representation, Semantic Gaussian, which is capable of preserving the scene’s 3D geometry, semantic labels associated with each Gaussian, and intricate texture details. Leveraging this novel representation of map, we directly predict the position of target object depicted in the goal image, thereby transforming IIN into a more tractable PointGoal Navigation task. Our proposed model achieves state-of-the-art performance, significantly enhancing SPL from 0.252 to 0.578. Furthermore, we analyze the error modes for our model and quantify the scope for improvement along two important dimensions (match and object grounding) in the future work.

### v3

@@ IV-D Efficiency Analysis
Efficiently locating the target object within the map is crucial for our method, which leverages renderings to match and ground the goal object. To reduce the computational cost of comparing all possible observations at each navigable viewpoint, we introduce Semantic Gaussians. This technique groups object instances under their corresponding semantic labels. This optimization significantly reduces the search space by limiting comparisons to several descriptive renderings of each instance. For example, in the scene CrMo8WxCyVb, the navigable area is quantified to be $54.03m^{2}$. This space can be discretized into approximately 54 squares, each with an area of $1m^{2}$. Positioned within each square, the agent can observe its surroundings from 12 distinct viewing angles, covering a full $360\deg$ with each angle spanning $30\deg$. Thus, the original search space for this single floor would consist of $12\times 54=648$ potential observations. By applying our grouping strategy based on semantic labeling, the search space is considerably narrowed. We count the number of different categories of object instances within the CrMo8WxCyVb scene, as demonstrated in Table IV. To locate a “chair” — assuming we render each object instance from three unique viewpoints — the resulting search space is reduced to merely $3\times 11=33$ observations. We render three observations of a single instance to ensure rendering quality, which is achievable only when the new viewpoint largely overlaps with the training views. This optimization yields a significant improvement in time efficiency.
@@
We also compare the runtime frame rates of different methods in a single Habitat environment using an NVIDIA GeForce RTX 3090 GPU and 10 CPU cores. As shown in Figure 7, our method maintains a high efficiency of over 20 FPS while achieving the highest SPL among the compared approaches. To achieve higher runtime speed, our GaussNav projects the Semantic Gaussian to obtain a 2D grid map. By utilizing the predicted target location from the “preprocessing”, we then employ Fast Marching Method (FMM) for path planning. The reason why our method outperforms various Modular Methods in terms of speed is that we do not rely on additional modules such as semantic segmentation [20, 22], local feature matching [20, 22], or switch module [22] during the navigation. This simplification enables GaussNav to operate more efficiently. We also provide a qualitative example of our GaussNav navigating to the target object instance, as depicted in Figure 9.
IV-E Error Analysis
The performance of the proposed model is still far from perfect. We would like to understand the error modes for future improvements. Our analysis identifies two sources of error: the first being the model’s inability to consistently match the target from instance renderings, and the second, inaccuracies in goal localization. To quantify the impact of these error sources, we conduct an evaluation of our model using a ground truth Match module and an accurate goal localization. The first one means agent can correctly recognize the target from candidate observations, and the second suggests agent is directly provided with the ground truth goal position. The variant equipped with a ground truth Match module (GaussNav w. GT Match in Table II) shows that Success can be enhanced by approximately 0.127 (rows 5 and 6 in Table II). Furthermore, when the model is augmented with both ground truth Match and Goal Localization, denoted as GaussNav w. GT Goal Localization, we observe an increase in Success from 0.850 to 0.946, as indicated in rows 6 and 7 in Table II. Improvements in the first error source may be achievable through the development of a more robust re-identification algorithm. As for the second source, a more precise Grounding strategy could yield better results. These insights not only highlight the model’s current shortcomings but also chart a course for subsequent refinement efforts.
IV-F Gaussian Construction Results
As illustrated in Figure 10, we provide visualization results of our Semantic Gaussian. These examples demonstrate the effectiveness of our Semantic Gaussian representation across a diverse range of scenarios. By presenting a more extensive collection of results, we aim to showcase the robustness and applicability of our approach in handling various scene complexities and object compositions.
@@ IV-F Gaussian Construction Results
In Figure 11, we present an quantitative evaluation of the rendering quality produced by our Semantic Gaussian Construction method on the HM3D validation dataset [14]. To align with the constraints of IIN task [19], we divide each scene within HM3D into separate floors and restrict the agent’s movement to within a single floor, as the IIN task [19] inherently ensures that both the agent’s starting location and the target’s position are on the same floor.
@@
To quantitatively analyze the results in Figure 11, we observe that the rendering results exhibit a bifurcated trend. For instance, in scenes with indices 29 and 30, the rendered images achieve a high PSNR of up to 40 and a near-zero depth rendering error. However, the rendering performance for scene 10 is suboptimal. We hypothesize that this polarized rendering quality across different scenes can be attributed to the discrepancy between the simulation and reality. This is evident in Figure 12, where some renderings from the HM3D dataset using Habitat simulator exhibit low fidelity, particularly in highly textured environments. High-quality reconstruction in such intricate settings is difficult, and utilizing suboptimal renderings as a basis for 3D environment reconstruction can further degrade the quality of the final output. In light of this, for scenes that are poorly reconstructed, we maintain consistency by using the original training views, rather than attempting to render novel views which would likely result in a diminished quality.
V Conclusion
@@
In this work, we introduce a modular approach for visual navigation, i.e., Gaussian Splatting for Visual Navigation (GaussNav). Previous map-based methods largely focus on building 2D BEV map, which cannot represent the 3D geometry and detailed features in a scene. To this end, we propose a novel map representation, Semantic Gaussian, which is capable of preserving the scene’s 3D geometry, semantic labels associated with each Gaussian, and intricate texture details. Leveraging this novel representation of map, we directly predict the position of target object depicted in the goal image, thereby transforming IIN into a more tractable PointGoal Navigation task. Our proposed framework achieves state-of-the-art performance, significantly enhancing SPL from 0.347 to 0.578. Furthermore, we analyze the error modes for our model and quantify the scope for improvement along two important dimensions (match and object grounding) in the future work.